If mayette is 6 years older than her sister beverly, and raul is 5 years older than mayette. It was known that thier total age is 41. How old is beverly?

Answer by rm29924(97)

Let Beverly's age = X
Let Mayette's age = X + 6
Let Raul's age = X + 6 + 5
All their ages added together = 41
So X + X + 6 + X + 6 + 5 = 41
combine like terms
3X + 17 = 41
subtract 17 from both sides
3X = 24
divide 3 from both sides to get the variable X by itself
X = 8
Beverly is 8 years old.